Prev Next

SW&IT Meeting Minutes 2020-04-20

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200406
LR20200406 - Bec, Matthieu 80.0
LR20200406 - Holmberg, Carl 72.0
LR20200406 - Edmeades, David 40.0
LR20200406 - Sola, Igor 80.0
LR20200406 - Hooper, Stephen 80.0
LR20200406 - Zhang, Yang 79.0
LR20200406 - Zhang, Xianyu 51.0

News from management

  • Board meeting Mon/Tue
  • IssueTrak initiative - Bucki's notes, IS as POC for future iterations
  • Covid19 Safety Protocols 2-sided pdf (CV's email, prepared by DC)

Previous Action Items

Owner Action Status
IS/DE create a new Luci simulator for ROWG purpose (+ work out resources available on the VM) existing OSA simulator will be fine
DE/XZ Xianyu's IDL personal license WIP - DE with Louise
MB SOUL/UAO follow-up CV in the loop, Julian LN/Argos, this Thursday meet with Al for LN, Argos TBD
IS dig-out how LUCI computes its WCS received document from Igor, next step Olga to follow up? Yang found some old emails

Open topics

  • Sabino power work Mon/Tue
    • Tue OVMS snapshot on UPS - Stephen will restart OVMs as needed
  • Steward power outage Fri/Sat - waiting to hear more
    • Expected to be same as last week - meaning no actions for this week.

Team reports

MB

  • Last week
    • OC, Management
    • student ad
    • ALTA discuss DMS, bug fix
    • Shark-VIS follow-up, USB3 single mode ok
    • found odd "sided forest" in OVMS/PSD vibration ?
    • AO meeting PB branch discussion
    • DMS API demo for Josh, updated refmem example
  • This week
    • board meeting
    • ALTA continued
    • AO / Arcetri meeting continued
    • phone interviews ?
    • cobbler / kickstart, ims f30
    • DMS weather loss, websockets, AdSec, etc.
    • Jira / Trak ?

DE

  • Steward power outage did not affect us or network; will announce next one as a matter of protocol but it's stated to have the same scope as Friday's
  • Started Raritan ticket; have some troubleshooting to do remotely
  • Everybody gets a Zoom room!
  • Decided to modify existing luci VM for current project; will delete the clone I did last week
  • Ordered new vibration monitoring laptop for Peter
  • Set meeting with Peter on Wednesday for finicky software install, supporting new vibration monitoring hardware
  • Ordered server room monitor hardware
  • Verified Stephen's suspicion re: PCI risers on Dell servers. Our "technical sales contact" had no idea that adapting PCIe to PCI was possible.
  • Veeam direct connect to VMHost1 easy, backups are proceeding. Still some snapshot stickiness; will work on clearing those this week
  • SMART failure on drive in Tucson Synology; will replace next time I have an office visit
  • Finding out from Louise about who is IDL license admin
  • Safety training
  • IssueTrak server accessibility issues seem to have been brief and have not reoccurred as far as I am aware.

YZ

  • Last week:
    • Continue IT#8021, LBC rotator trailing issues.
    • DS9 XPA study (for modsAlign refractoring).
    • AO-Cravat work on refractoring Strehl Python code, plus recomputing Strehl values on Al's current OSCO data.
    • TMS Mathematica code follow-up (email TMS list/Andrew for questions).
    • Discussed with Olga on setting up local modsAlign test (she sent several test files and instructions).
    • Updating TCS/IIF interface document (unofficial version).
  • This week:
    • Continue IT#8021, LBC rotator trailing issues. Current plan is to communicate with Olga and support her needs.
    • Finish updating TCS/IIF interface document (unofficial version).
    • Set up local modsAlign testing and experiment on DS9 XPA communication and interaction.
    • Follow up on previous AO-Cravat work (also github setup).
    • Follow up on TMS Mathematica->Python work.
    • Keep conversation on MDOS WCS keywords going with Olga.
    • Safety training.

CH

  • Last week:
    • DMS: revisited approach to storing data for repurposed iMonnit sensors.
    • indi-core: completed rpm packaging.
  • This week:
    • DMS: implement repurposed sensor data storage, add newly installed sensors.
    • indi-core: complete testing of software as packaged.

XZ

  • Last week
    • Actuators status analysis
      • transfer data from NAS to my laptop
      • add code for configration detection
      • modify code for dataset detection method
      • test and verify with data
    • PB SW
      • AO SW versions merge meeting and preparation
      • AO SW and new 4D SW preparation and meeting
      • MG shared document setup test
      • PB SW plan meeting
      • ao-confcalib dicussion
      • install and test MATLAB for PB lab test
      • copy MG firmware and MATLAB packages to my laptop
    • Meetings
  • This week
    • PB action items
    • PB SW/HW

SH

  • ovirt mostly setup
    • hyper-converged (need to grok a little more)
    • moved over three VMs (also upgraded to Fedora 31)
  • puchased monitor for ROWG
  • laptop up mountain for Pat
  • SSL for LBTC
  • Mountain power down
  • Mountain MTU
  • Core-B (amber light)
  • meetings

next:

  • Tucson VPN
  • 507 upgrades and new machine for ROWG.
  • AllSky

IS

Last week:
  • PIT
    • Fixed declination issue in italian pit
    • working on submission issues on Mac
    • PDF table mismatch in column for the general pit.
    • Made new release INAF PIT 2.1004.0 test release
    • Worked on PIT admin tool
  • OT
    • working on bug fixes, hopefully a test release will go out soon.

This week:
  • PIT
    • If everything goes well do final release for INAF
    • Keep working on the PIT admin tool backend.
  • OT
    • work with Andrew on current test OT release issues
  • LUCI
    • install latest GEIRS in simulation environment
    • Make sure OSA-luci is uptodate.
    • Look into LUCI WebIO device configuration.
Topic revision: r15 - 20 Apr 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback